# Infinite graph product of groups I: Geometry of the extension graph

### v2

@@ 3.1 Definition and basic properties of the extension graph of graph product
(1) Without loss of generality, we assume $a\in V(\Gamma)$. Let $g\in\Gamma\mathcal{G}$ satisfy $b=g.v(b)$. By $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)\neq\{1\}$, there exist $k\in\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\setminus\{1\}$ and $h\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v(b))\setminus\{1\}$ such that $k=ghg^{-1}$. By Corollary 3.6, we have $\emptyset\neq\mathrm{supp}(k)\subset\mathrm{St}_{\Gamma}(a)$ and $\emptyset\neq\mathrm{supp}(h)\subset\mathrm{St}_{\Gamma}(v(b))$. By Lemma 3.15, we have $\mathrm{supp}(h)\cap\mathrm{supp}(ghg^{-1})\neq\emptyset$ and there exist $k_{1},k_{2},k_{3}\in\Gamma\mathcal{G}$ and $w\in\mathrm{St}_{\Gamma}(v(b))$ such that $g=k_{1}k_{2}k_{3}$, $\mathrm{supp}(k_{1})\subset\mathrm{supp}(ghg^{-1})=\mathrm{supp}(k)\subset\mathrm{St}_{\Gamma}(a)$, $\mathrm{supp}(k_{2})\subset\mathrm{St}_{\Gamma}(w)$, and $\mathrm{supp}(k_{3})\subset\mathrm{St}_{\Gamma}(v(b))$. This implies $b=k_{1}k_{2}k_{3}.v(b)=k_{1}k_{2}.v(b)$, $k_{1}k_{2}.w=k_{1}.w$, and $a=k_{1}.a$. By $\emptyset\neq\mathrm{supp}(h)\cap\mathrm{supp}(ghg^{-1})\subset\mathrm{St}_{\Gamma}(v(b))\cap\mathrm{St}_{\Gamma}(a)$, we also have $d_{\Gamma}(a,v(b))\leq 2$. Hence, we have
$d_{\Gamma^{e}}(a,b)=d_{\Gamma^{e}}(a,k_{1}k_{2}.v(b))\leq d_{\Gamma^{e}}(k_{1}.a,k_{1}.w)+d_{\Gamma^{e}}(k_{1}k_{2}.w,k_{1}k_{2}.v(b))\leq 3+1=4.$
(2) Let $h\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(v)g^{-1}$ with $h\neq 1$. There exists $h^{\prime}\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\setminus\{1\}$ such that $h=gh^{\prime}g^{-1}$. By Corollary 3.6, we have $\emptyset\neq\mathrm{supp}(h^{\prime})\subset\mathrm{St}_{\Gamma}(v)$. By this and Lemma 3.15, there exist $k_{1},k_{2},k_{3}\in\Gamma\mathcal{G}$ and $w\in\mathrm{St}_{\Gamma}(v)$ such that $g=k_{1}k_{2}k_{3}$, $\mathrm{supp}(k_{1})\subset\mathrm{supp}(gh^{\prime}g^{-1})=\mathrm{supp}(h)\subset\mathrm{St}_{\Gamma}(v)$, $\mathrm{supp}(k_{2})\subset\mathrm{St}_{\Gamma}(w)$, and $\mathrm{supp}(k_{3})\subset\mathrm{St}_{\Gamma}(v)$. We can assume $\|k_{1}k_{2}k_{3}\|=\|k_{1}\|+\|k_{2}\|+\|k_{3}\|$ by canceling syllables in normal forms of $k_{1}$, $k_{2}$, and $k_{3}$. By $g\notin\mathrm{Stab}_{\Gamma\mathcal{G}}(v)$ and $\mathrm{supp}(k_{1})\cup\mathrm{supp}(k_{3})\subset\mathrm{St}_{\Gamma}(v)$, we have $\mathrm{supp}(k_{2})\setminus\mathrm{St}_{\Gamma}(v)\neq\emptyset$. Hence, there exists a geodesic decomposition $k_{1}=w_{1}sw_{2}$ (i.e. $\|k_{2}\|=\|w_{1}\|+\|s\|+\|w_{2}\|$), where $w_{1},w_{2}\in\Gamma\mathcal{G}$, $s\in G_{z}$, and $z\in V(\Gamma)$, such that $z\in\mathrm{St}_{\Gamma}(w)\setminus\mathrm{St}_{\Gamma}(v)$ and $\mathrm{supp}(w_{2})\subset\mathrm{St}_{\Gamma}(v)$. By $\mathrm{supp}(gh^{\prime}g^{-1})=\mathrm{supp}(h)\subset\mathrm{St}_{\Gamma}(v)$, the syllable $s$ must cancel out with $s^{-1}$ in $k_{2}^{-1}=w_{2}^{-1}s^{-1}w_{1}^{-1}$ to get a normal form of $gh^{\prime}g^{-1}$. Hence, we can see $\mathrm{supp}(w_{2}k_{3}h^{\prime}k_{3}^{-1}w_{2}^{-1})\subset\mathrm{Lk}_{\Gamma}(z)$. By $\mathrm{girth}(\Gamma)>4$, we have $\mathrm{Lk}_{\Gamma}(z)\cap\mathrm{St}_{\Gamma}(v)=\{w\}$. Hence, $w_{2}k_{3}h^{\prime}k_{3}^{-1}w_{2}^{-1}\in G_{w}$. Define $g_{1}\in\Gamma\mathcal{G}$ by $g_{1}=k_{1}w_{1}s$, then we have $h\in g_{1}G_{w}g_{1}^{-1}$. Since we can take $g_{1}$ independent of $h$, this implies $\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(v)g^{-1}\subset g_{1}G_{w}g_{1}^{-1}$.
(3) We assume $\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(w)g^{-1}\neq\{1\}$. Otherwise the statement is trivial. Let $h\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(w)g^{-1}$ with $h\neq 1$. There exists $h^{\prime}\in\mathrm{Stab}_{\Gamma\mathcal{G}}(w)\setminus\{1\}$ such that $h=gh^{\prime}g^{-1}$. By Corollary 3.6, we have $\emptyset\neq\mathrm{supp}(gh^{\prime}g^{-1})=\mathrm{supp}(h)\subset\mathrm{St}_{\Gamma}(v)$ and $\emptyset\neq\mathrm{supp}(h^{\prime})\subset\mathrm{St}_{\Gamma}(w)$. Hence, by Lemma 3.15 we have $\mathrm{supp}(h^{\prime})\cap\mathrm{supp}(gh^{\prime}g^{-1})\neq\emptyset$ and there exist $k_{1},k_{2},k_{3}\in\Gamma\mathcal{G}$ and $z\in\mathrm{St}_{\Gamma}(w)$ such that $g=k_{1}k_{2}k_{3}$, $\mathrm{supp}(k_{1})\subset\mathrm{supp}(gh^{\prime}g^{-1})\subset\mathrm{St}_{\Gamma}(v)$, $\mathrm{supp}(k_{2})\subset\mathrm{St}_{\Gamma}(z)$, and $\mathrm{supp}(k_{3})\subset\mathrm{St}_{\Gamma}(w)$. By $\emptyset\neq\mathrm{supp}(h^{\prime})\cap\mathrm{supp}(gh^{\prime}g^{-1})\subset\mathrm{St}_{\Gamma}(w)\cap\mathrm{St}_{\Gamma}(v)$, we have $d_{\Gamma}(v,w)\leq 2$. We’ll discuss two cases, (i) when $d_{\Gamma}(v,w)=2$ and (ii) when $d_{\Gamma}(v,w)=1$.
@@
(i) When $d_{\Gamma}(v,w)=2$, by $\mathrm{girth}(\Gamma)>4$, there exists $u\in V(\Gamma)$ such that $\{u\}=\mathrm{St}_{\Gamma}(v)\cap\mathrm{St}_{\Gamma}(w)$. By $\mathrm{girth}(\Gamma)>5$ and $z\in\mathrm{St}_{\Gamma}(w)$, we have $\mathrm{St}_{\Gamma}(v)\cap(\mathrm{St}_{\Gamma}(w)\cup\mathrm{St}_{\Gamma}(z))\subset\{v,u\}$. Note $\mathrm{supp}(k_{2}k_{3}h^{\prime}k_{3}^{-1}k_{2}^{-1})\subset\mathrm{St}_{\Gamma}(w)\cup\mathrm{St}_{\Gamma}(z)$. Also, $\mathrm{supp}(k_{2}k_{3}h^{\prime}k_{3}^{-1}k_{2}^{-1})=\mathrm{supp}(k_{1}^{-1}hk_{1})\subset\mathrm{St}_{\Gamma}(v)$. Hence, we have $\mathrm{supp}(k_{2}k_{3}h^{\prime}k_{3}^{-1}k_{2}^{-1})\subset\{u,v\}$. This implies $h\in k_{1}(G_{v}\times G_{u})k_{1}^{-1}$ by Remark 3.12 (1). Hence, we have $|\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(w)g^{-1}|\leq|G_{v}||G_{u}|<\infty$ since we can take $k_{1}$ independent of $h$.
(ii) When $d_{\Gamma}(v,w)=1$. If $z\neq v$, then $\mathrm{girth}(\Gamma)>4$ implies $\mathrm{St}_{\Gamma}(v)\cap(\mathrm{St}_{\Gamma}(w)\cup\mathrm{St}_{\Gamma}(z))\subset\{v,w\}$, hence we can see $h\in k_{1}(G_{v}\times G_{w})k_{1}^{-1}$ in the same way as case (i). If $z=v$, then $\mathrm{supp}(k_{1}k_{2})\subset\mathrm{St}_{\Gamma}(v)$. Note $\mathrm{supp}(k_{3}h^{\prime}k_{3}^{-1})\subset\mathrm{St}_{\Gamma}(w)$. Also, $\mathrm{supp}(k_{3}h^{\prime}k_{3}^{-1})=\mathrm{supp}(k_{2}^{-1}k_{1}^{-1}hk_{1}k_{2})\subset\mathrm{St}_{\Gamma}(v)$. Hence, $\mathrm{supp}(k_{3}h^{\prime}k_{3}^{-1})\subset\mathrm{St}_{\Gamma}(v)\cap\mathrm{St}_{\Gamma}(w)=\{v,w\}$. This implies $h\in k_{1}k_{2}(G_{v}\times G_{w})k_{2}^{-1}k_{1}^{-1}$ by Remark 3.12 (1). Hence, $|\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap g\mathrm{Stab}_{\Gamma\mathcal{G}}(w)g^{-1}|\leq|G_{v}||G_{w}|<\infty$ since we can take $k_{1}$ and $k_{2}$ independent of $h$. ∎

### v3

@@ 3.1 Definition and basic properties of the extension graph of graph product
(1) We first show that $d_{\Gamma^{e}}(a,b)>2$ implies $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)=\{1\}$. To show the contraposition, suppose that $a,b\in V(\Gamma^{e})$ satisfy $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)\neq\{1\}$. Without loss of generality, we assume $a\in V(\Gamma)$. Let $g\in\Gamma\mathcal{G}$ satisfy $b=g.v(b)$. By $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)\neq\{1\}$, there exist $k\in\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\setminus\{1\}$ and $h\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v(b))\setminus\{1\}$ such that $k=ghg^{-1}$. By Corollary 3.6, we have $\emptyset\neq\mathrm{supp}(k)\subset\mathrm{St}_{\Gamma}(a)$ and $\emptyset\neq\mathrm{supp}(h)\subset\mathrm{St}_{\Gamma}(v(b))$. By Lemma 3.15, we have $\mathrm{supp}(h)\cap\mathrm{supp}(ghg^{-1})\neq\emptyset$ and there exist $k_{1},k_{2},k_{3}\in\Gamma\mathcal{G}$ and $w\in\mathrm{St}_{\Gamma}(v(b))$ such that $g=k_{1}k_{2}k_{3}$, $\mathrm{supp}(k_{1})\subset\mathrm{supp}(ghg^{-1})=\mathrm{supp}(k)\subset\mathrm{St}_{\Gamma}(a)$, $\mathrm{supp}(k_{2})\subset\mathrm{St}_{\Gamma}(w)$, and $\mathrm{supp}(k_{3})\subset\mathrm{St}_{\Gamma}(v(b))$. Here, we retake $k_{2}$ and $k_{3}$ so that $\|k_{2}\|$ is the minimum among all $\|k_{2}^{\prime}\|$, where $k_{2}^{\prime},k_{3}^{\prime}\in\Gamma\mathcal{G}$ satisfy $k_{2}k_{3}=k_{2}^{\prime}k_{3}^{\prime}$, $\mathrm{supp}(k_{2}^{\prime})\subset\mathrm{St}_{\Gamma}(w)$, and $\mathrm{supp}(k_{3}^{\prime})\subset\mathrm{St}_{\Gamma}(v(b))$. In particular, $\|k_{2}(k_{3}hk_{3}^{-1})\|=\|k_{2}\|+\|k_{3}hk_{3}^{-1}\|$ and $\|(k_{3}hk_{3}^{-1})k_{2}^{-1}\|=\|k_{3}hk_{3}^{-1}\|+\|k_{2}^{-1}\|$ hold. Also, by $\emptyset\neq\mathrm{supp}(h)\cap\mathrm{supp}(ghg^{-1})\subset\mathrm{St}_{\Gamma}(v(b))\cap\mathrm{St}_{\Gamma}(a)$, we have $d_{\Gamma}(a,v(b))\leq 2$. We will discuss two cases (i) and (ii).
@@
(i) When $\|k_{2}(k_{3}hk_{3}^{-1})k_{2}^{-1}\|<\|k_{2}\|+\|k_{3}hk_{3}^{-1}\|+\|k_{2}^{-1}\|$, by $\|k_{2}(k_{3}hk_{3}^{-1})\|=\|k_{2}\|+\|k_{3}hk_{3}^{-1}\|$ and $\|(k_{3}hk_{3}^{-1})k_{2}^{-1}\|=\|k_{3}hk_{3}^{-1}\|+\|k_{2}^{-1}\|$, there exist a syllable $s$ of $k_{2}$ and a syllable $s^{\prime}$ of $k_{2}^{-1}$ that cancel in $k_{2}(k_{3}hk_{3}^{-1})k_{2}^{-1}$. Hence, we have $\mathrm{supp}(k_{3}hk_{3}^{-1})\subset\mathrm{Lk}_{\Gamma}(\mathrm{supp}(s))$ and also $\mathrm{supp}(s)\notin\mathrm{St}_{\Gamma}(v(b))$ by the minimality of $\|k_{2}\|$ (see Remark 2.16). By $\mathrm{supp}(k_{3}hk_{3}^{-1})\subset\mathrm{Lk}_{\Gamma}(\mathrm{supp}(s))\cap\mathrm{St}_{\Gamma}(v(b))$, $\mathrm{supp}(s)\in\mathrm{St}_{\Gamma}(w)\setminus\mathrm{St}_{\Gamma}(v(b))$, and $\mathrm{girth}(\Gamma)>4$, we have $\mathrm{supp}(k_{3}hk_{3}^{-1})=\{w\}$. By this, $k=(k_{1}k_{2})(k_{3}hk_{3}^{-1})(k_{1}k_{2})^{-1}$, and Lemma 3.4, we have $w\in\mathrm{supp}(k)\subset\mathrm{St}_{\Gamma}(a)$. Thus, by $b=k_{1}k_{2}k_{3}.v(b)=k_{1}k_{2}.v(b)$, $k_{1}k_{2}.w=k_{1}.w$, and $a=k_{1}.a$, we have $d_{\Gamma^{e}}(a,b)=d_{\Gamma^{e}}(a,k_{1}k_{2}.v(b))\leq d_{\Gamma^{e}}(k_{1}.a,k_{1}.w)+d_{\Gamma^{e}}(k_{1}k_{2}.w,k_{1}k_{2}.v(b))\leq 1+1=2$.
(ii) When $\|k_{2}(k_{3}hk_{3}^{-1})k_{2}^{-1}\|=\|k_{2}\|+\|k_{3}hk_{3}^{-1}\|+\|k_{2}^{-1}\|$, we have $\mathrm{supp}(k_{2})\cup\mathrm{supp}(k_{3}hk_{3}^{-1})=\mathrm{supp}(k_{2}(k_{3}hk_{3}^{-1})k_{2}^{-1})=\mathrm{supp}(k_{1}^{-1}kk_{1})\subset\mathrm{St}_{\Gamma}(a)$. This implies $a=k_{1}k_{2}.a$. By this and $b=k_{1}k_{2}k_{3}.v(b)=k_{1}k_{2}.v(b)$, we have $d_{\Gamma^{e}}(a,b)=d_{\Gamma^{e}}(k_{1}k_{2}.a,k_{1}k_{2}.v(b))=d_{\Gamma^{e}}(a,v(b))\leq 2$.
By (i) and (ii), we have shown $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)\neq\{1\}$ implies $d_{\Gamma^{e}}(a,b)\leq 2$.
Next, suppose $d_{\Gamma^{e}}(a,b)=2$ and let $c\in V(\Gamma^{e})$ satisfy $\{(a,c),(c,b)\}\subset E(\Gamma^{e})$. By Lemma 3.8, we may assume $\{a,c\}\subset V(\Gamma)$ without loss of generality. Let $g\in\Gamma\mathcal{G}$ satisfy $(c,b)=g.(v(c),v(b))$ (such $g$ exists by Lemma 3.8). By $c\in V(\Gamma)$ and Corollary 3.7, we have $v(c)=c$, hence $g\in\mathrm{Stab}_{\Gamma\mathcal{G}}(c)$.
Let $k\in\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)$. By $\mathrm{Stab}_{\Gamma\mathcal{G}}(b)=g\mathrm{Stab}_{\Gamma\mathcal{G}}(v(b))g^{-1}$, there exists $h\in\mathrm{Stab}_{\Gamma\mathcal{G}}(v(b))$ such that $k=ghg^{-1}$. Note $\mathrm{supp}(k)\subset\mathrm{St}_{\Gamma}(a)$, $\mathrm{supp}(g)\subset\mathrm{St}_{\Gamma}(c)$, and $\mathrm{supp}(h)\subset\mathrm{St}_{\Gamma}(v(b))$ by Corollary 3.6. We will discuss two cases (i’) and (ii’).
(i’) When $a\neq v(b)$, we have $\mathrm{St}_{\Gamma}(a)\cap\big(\mathrm{St}_{\Gamma}(c)\cup\mathrm{St}_{\Gamma}(v(b))\big)=\{a,c\}$ and $\mathrm{St}_{\Gamma}(c)\cap\mathrm{St}_{\Gamma}(v(b))=\{c,v(b)\}$ by $\mathrm{girth}(\Gamma)>4$. Hence, we can see $\mathrm{supp}(k)\subset\{a,c\}$ and $\mathrm{supp}(h)\subset\{c,v(b)\}$ by comparing the support of the syllables on each side of the equality $k=ghg^{-1}$. In particular, $\{k,g,h\}\subset\mathrm{Stab}_{\Gamma\mathcal{G}}(c)=G_{c}\times\big(*_{v\in\mathrm{Lk}_{\Gamma}(c)}G_{v}\big)$ (note $\mathrm{girth}(\Gamma)>4$). Hence, we have $\mathrm{supp}(h)\subset\{c\}$, which implies $\mathrm{supp}(k)\subset\{c\}$. Indeed, $v(b)\in\mathrm{supp}(h)$ would imply $v(b)\in\mathrm{supp}(k)$, which contradicts $\mathrm{supp}(k)\subset\{a,c\}$.
(ii’) When $a=v(b)$, by $a\neq b~(=g.a)$, there exists a syllable $s$ of $g$ such that $\mathrm{supp}(s)\notin\{a,c\}$. Since $s$ must cancel in $ghg^{-1}$ by $\mathrm{supp}(s)\cap\mathrm{supp}(k)\subset\mathrm{supp}(s)\cap\mathrm{St}_{\Gamma}(a)=\emptyset$ but cannot cancel with any syllable of $h$ by $\mathrm{supp}(s)\cap\mathrm{supp}(h)\subset\mathrm{supp}(s)\cap\mathrm{St}_{\Gamma}(a)=\emptyset$, we can see $\mathrm{supp}(h)\subset\{a,c\}$. In particular, $\{g,h\}\subset\mathrm{Stab}_{\Gamma\mathcal{G}}(c)=G_{c}\times\big(*_{v\in\mathrm{Lk}_{\Gamma}(c)}G_{v}\big)$. Hence, by $\mathrm{supp}(s)\notin\{a,c\}$ and $\mathrm{supp}(k)=\mathrm{supp}(ghg^{-1})\subset\mathrm{St}_{\Gamma}(a)\cap\mathrm{St}_{\Gamma}(c)=\{a,c\}$, we can see $\mathrm{supp}(h)\subset\{c\}$, which implies $\mathrm{supp}(k)\subset\{c\}$.
By (i’) and (ii’), we have shown $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)\subset G_{c}$. The converse inclusion is straightforward. By a similar argument on the support of syllables, we can also show $\mathrm{girth}(\Gamma^{e})>4$ using $\mathrm{girth}(\Gamma)>4$. Hence, we have $\{G_{c}\}=\mathrm{St}_{\Gamma^{e}}(a)\cap\mathrm{St}_{\Gamma^{e}}(b)$.
Finally, when $d_{\Gamma^{e}}(a,b)=1$, we can see $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(b)=a\times b$ by Lemma 3.8 and Remark 3.12 (1). Here, note $\mathrm{St}_{\Gamma^{e}}(a)\cap\mathrm{St}_{\Gamma^{e}}(b)=\{a,b\}$ by $\mathrm{girth}(\Gamma^{e})>4$. The case $d_{\Gamma^{e}}(a,b)=0$ follows from Corollary 3.6. Indeed, assume $a\in V(\Gamma)$ without loss of generality, then we have $\mathrm{Stab}_{\Gamma\mathcal{G}}(a)=\langle G_{v}\mid v\in\mathrm{St}_{\Gamma}(a)\rangle=\langle H\mid H\in\mathrm{St}_{\Gamma^{e}}(a)\rangle$. Here, the second equality holds since for any $g\in\Gamma\mathcal{G}$ and $v\in V(\Gamma)$ satisfying $gG_{v}g^{-1}\in\mathrm{Lk}_{\Gamma^{e}}(a)$, we have $g\in\mathrm{Stab}_{\Gamma\mathcal{G}}(a)$ and $v\in\mathrm{Lk}_{\Gamma}(a)$ by Corollary 3.7 and Lemma 3.8.
(2) By $g\in\Gamma\mathcal{G}\setminus\mathrm{Stab}_{\Gamma\mathcal{G}}(v)$ and Lemma 3.8, we have $d_{\Gamma^{e}}(v,g.v)\geq 2$. By the proof of Corollary 3.16 (1), the group $\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(g.v)$ is $\{1\}$ if $d_{\Gamma^{e}}(v,g.v)>2$ and $c$ if $d_{\Gamma^{e}}(v,g.v)=2$ where $\{c\}=\mathrm{St}_{\Gamma^{c}}(v)\cap\mathrm{St}_{\Gamma^{c}}(g.v)$.
(3) Note $g\mathrm{Stab}_{\Gamma\mathcal{G}}(w)g^{-1}=\mathrm{Stab}_{\Gamma\mathcal{G}}(g.w)$. By $v\neq w$ and Corollary 3.7, we have $v\neq g.w$. Hence, by the proof of Corollary 3.16 (1), the group $\mathrm{Stab}_{\Gamma\mathcal{G}}(v)\cap\mathrm{Stab}_{\Gamma\mathcal{G}}(g.w)$ is $\{1\}$ if $d_{\Gamma^{e}}(v,g.w)>2$, $c$ if $d_{\Gamma^{e}}(v,g.w)=2$ where $\{c\}=\mathrm{St}_{\Gamma^{c}}(v)\cap\mathrm{St}_{\Gamma^{c}}(g.w)$, and $v\times g.w$ if $d_{\Gamma^{e}}(v,g.w)=1$. Here, $v$, $c$, and $g.w$ are finite subgroups of $\Gamma\mathcal{G}$ by the assumption on $\mathcal{G}$. ∎